Circular No. 6526

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 960720 C. Luginbuhl and F. Vrba, U.S. Naval Observatory, Flagstaff Station;

R. Hudec, P. Pravec, and L. Šarounová, Astronomical Institute, Ondřejov; J. Greiner, Max-Planck-Institut für Extraterrestrische Physik, Garching; and G. Tovmassian, Instituto de Astronomía, Universidad Nacional Autonoma de México (UNAM), report: “Broadband BVR CCD frames taken at the Ondřejov 0.65-m telescope have revealed a single very blue object within the 10′-radius Beppo-SAX error circle for GRB 960720 (cf. IAUC 6467) at α = 17^{h}29^{m}24^{s}.73, δ = +49^{o}06′22′′.0 (equinox 2000.0); these images do not reach deep enough to detect the fainter objects seen on the POSS images. Broadband UBVI photometric observations of this object obtained at the Flagstaff Station 1.0-m telescope on Oct. 9 and 10 UT show this object to have colors typical of quasistellar objects (QSOs), with V = 18.25 ± 0.01, U − B = −0.31 ± 0.08, B − V = +1.21 ± 0.05, V − I = +1.36 ± 0.02. An optical spectrum (range 400–670 nm) obtained on Oct. 19 with the UNAM 2.15-m telescope at Cananea confirms the QSO (which was outside the 1′- radius error circle of the quiescent x-ray source discussed on IAUC 6481), showing a strong, broad emission line at 418 nm and two fainter broad lines at 442 and 515 nm, which we interpret as the redshifted lines of C IV] at 155 nm, He II at 164 nm, and C III] at 191 nm, resulting in a redshift of

z = 1.7 (see http://www.rosat.mpe-garching.mpg.de/~jcg/grb.html).

This QSO is the likely counterpart of RX J1729.4+4906, mentioned (but not specificially designated) on IAUC 6487. From the Boyle et al. (1988, MNRAS 235, 935) QSO statistics, the likelihood of finding a QSO of this brightness or brighter within the Beppo-SAX localization by chance alone is 0.08. The Flagstaff Station observations also covered the other seven ROSAT sources within the Beppo-SAX localization described on IAUC 6487. For the Seyfert galaxy reported within the RX J1729.5+4904 error circle, our measures indicate a possible variation at U of 0.50 ± 0.16 mag between the two successive nights with mean photometric values of V = 20.39 ± 0.05, U − B = −0.59 ± 0.25, B − V = +0.28 ± 0.05, V − I = +0.40 ± 0.17. Follow-up observations of this Seyfert object are encouraged to confirm and characterize any variation.”

COMET 22P/KOPFF G. Cremonese, Astronomical Observatory, Padua, informs us that the

account of this comet’s splitting is in error and that the part of the report on IAUC 6521 discussing the splitting should be ignored.

1996 December 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green